OpenStudy (anonymous):

Estimate the sum of 9.327 + 5.132 to one decimal place.

OpenStudy (mathmate):

For an estimation, you could truncate to 1 decimal place and add: 9.3+5.1=14.4 To get a better estimate, add the second decimal place, if it is 5 or more, then round up the estimate. Second decimal place = 2+3=5 >=5, so round up 14.4 to 14.5.

OpenStudy (anonymous):